For each of the following variables, identify the type of variable (categorical vs. numeric).
(I) Number of auto insurance claims in a month
(II) Film genre (e.g. comedy, horror, drama, etc)

Question 1 options:

1)

(I) Categorical , and (II) Categorical

2)

There is no correct match.

3)

(I) Categorical , and (II) Numeric

4)

(I) Numeric , and (II) Numeric

5)

(I) Numeric , and (II) Categorical

Question 2 (1 point)

Saved

For each of the following variables, identify the level of measurement (nominal, ordinal, interval, or ratio).
(I) Satisfaction level on a survey (e.g. very dissatisfied, satisfied, very satisfied, etc.)
(II) Grade point average (e.g. 1.0, 2.0, 3.0 , 4.0)

Question 2 options:

1)

(I) Ratio , and (II) Nominal

2)

(I) Ordinal , and (II) Interval

3)

(I) Ordinal , and (II) Ratio

4)

(I) Ratio , and (II) Interval

5)

(I) Nominal , and (II) Nominal

Question 3 (1 point)

Saved

Choose the description which best describes the shape of this boxplot.
50.0 50.5 51.0 51.5 52.0 52.5

Question 3 options:

1)

Bimodal

2)

Symmetric

3)

Right-skewed

4)

Left-skewed

5)

Contaminated

Question 4 (1 point)

Choose the description which best describes the shape of this boxplot.
50 52 54 56

Question 4 options:

1)

Bimodal

2)

Left-skewed

3)

Right-skewed

4)

Symmetric

5)

Contaminated

Question 5 (1 point)

The five-number summary for a dataset on Hourly Unit Sales is given in the table below. Using this information, calculate the percent of values between 1 and 8.
Min Q1 Median Q3 Max 1 8 20 26 29

Question 5 options:

1)

0%

2)

100%

3)

50%

4)

75%

5)

25%

Question 6 (1 point)

The lifespans (in years) of ten beagles were 10; 7; 12; 10; 12; 12; 10; 6; 11; 10. Calculate the range of the dataset.

Question 6 options:

1)

9

2)

1.8

3)

2.05

4)

6

5)

10

Question 7 (1 point)

The site traffic for a business's website was monitored for 10 days, showing daily traffic of 15,267; 14,850; 15,854; 15,169; 14,454; 15,093; 15,334; 14,207; 13,908; 14,763. Calculate the coefficient of variation of the dataset.

Question 7 options:

1)

15,469.77

2)

1,159.74

3)

0.04

4)

-14,310.03

5)

8,634,215

Question 8 (1 point)

A new marketing campaign is being released by Pepsico to promote their Rockstar engergy drink line this summer and is expecting an increase in sales volume. Currently daily sales average $1,587.38 with a standard deviation of $16.315. Daily sales figures are compiled in a dataset and to forecast the change, 31.48 is added to every observation in the dataset. What is the new standard deviation for daily sales?

Question 8 options:

1)

47.8

2)

513.6

3)

1,618.86

4)

49,970.72

5)

16.32

Question 9 (1 point)

A breeder is trying to forcast the number of dogs they will need to house each year for the next five years. As a general rule, the average and standard deviation of the breeding pairs on hand are 9 and 2.2 respectively. To predict the number of puppies born in each of the next five years, every observation in the dataset is multiplied by 1.4. What will the new standard deviation be?

Question 9 options:

1)

3.08

2)

2.2

3)

12.6

4)

3.6

5)

10.4

Question 10 (1 point)

Suppose that the middle 68% of heights of adult males fall between 61.4 and 65.96. Give an approximate estimate of the standard deviation of heights. Assume that height has a normal distribution.

Question 10 options:

1)

2.28

2)

4.56

3)

63.68

4)

-2.28

5)

1.14

Question 11 (1 point)

The average number of typical text pages a standard print cartridge produces is 305.9 pages based on 9 printers. After the 10th printer is tested, the average is 306.2 pages. How many pages did the last cartridge produce?

Question 11 options:

1)

0

2)

0

3)

306

4)

309

5)

306

Question 12 (1 point)

Suppose that the scores on a statistics final exam have a mean of 87.79. If the scores are symmetric (e.g. in the plot below), will the median be larger, smaller, or approximately equal to 87.79?
Final Exam Score

Question 12 options:

1)

The median is greater than 87.79.

2)

The median is less than 87.79.

3)

We cannot calculate the median in a symmetric distribution.

4)

The median is approximately equal to 87.79.

5)

We do not have enough information to determine if the mean or median is larger.
Answer #1

1)\(I) Numeric , and (II) Categorical

2)

(I) Ordinal , and (II) Interval

3) Left-skewed

4) Symmetric

5)25%

6)range =max-min=12-6=6

7)coefficient of variation =(std deviation/mean)*100 =0.04

8) 16.32

9) new standard deviation be =1.4*2.2=3.08

10) approximate estimate of the standard deviation of heights =(65.96-61.4)/2=2.28

11) pages did the last cartridge produce =306.2*10-9*305.9=309

12) The median is approximately equal to 87.79.